# Amortizing Perpetual Options

### v1

@@ Abstract
In this work, we introduce amortizing perpetual options (AmPOs), a fungible variant of continuous-installment options suitable for exchange-based trading. Traditional installment options lapse when holders cease their payments, destroying fungibility across units of notional. AmPOs replace explicit installment payments and the need for lapsing logic with an implicit payment scheme via a deterministic decay in the claimable notional. This amortization ensures all units evolve identically, preserving fungibility. Under the Black-Scholes framework, AmPO valuation can be reduced to an equivalent vanilla perpetual American option on a dividend-paying asset. In this way, analytical expressions are possible for the exercise boundaries and risk-neutral valuations for calls and puts. These formulas and relations allow us to derive the Greeks and study comparative statics with respect to the amortization rate. Illustrative numerical case studies demonstrate how the amortization rate shapes option behavior and reveal the resulting tradeoffs in the effective volatility sensitivity.
@@ 1 Introduction
Perpetual contracts are experiencing a renaissance due to their adoption in decentralized finance and cryptocurrency markets. Though introduced in [17] for, e.g., a market on real estate indices, perpetual futures on cryptocurrencies are liquidly traded on a number of markets such as Hyperliquid and Binance. These contracts readily allow traders to purchase leveraged positions or to short assets which, otherwise, are technologically incapable of such a position; we refer the interested reader to, e.g., [1, 3]. However, perpetual futures only provide users with linear payoffs. Perpetual options, which replicate elements of the funding rate structure of perpetual futures, have recently been introduced in decentralized systems. We highlight everlasting options [19] and Panoptic [14] as specific perpetual option variants.
However, though perpetual futures have found wide usage in cryptocurrencies, perpetual options have not gained the same market share. Notably, the stochastic funding rates used in everlasting options and at Panoptic can be viewed as types of installment options. An installment option is one in which cash flows are exchanged between the long and short positions [8, 4]. Traditionally, these cash flows exist solely from the long to the short holders. Such options, however, only trade over-the-counter due to the nature of these contracts [4]. Specifically, installment options grant the holder a dual optionality: exercise or lapse. Because the lapsing logic comes from failure to make the installment payment in full, the notional units of these contracts are not identical and, therefore, cannot trade on an exchange. For the purposes of this work, we specifically highlight continuous-installment (CI) options which require a continuous stream of payments from the option holder to the option underwriter in order to keep the contract alive [12, 13].
The remainder of this paper is organized as follows. In Section 2, we introduce an exchange-tradable variant of CI options; this is accomplished by replacing the explicit installment payments with implicit amortization of the options notional. In Section 3, under the Black-Scholes framework, we study the risk-neutral valuation and Greeks by determining an equivalence with vanilla perpetual American options for a dividend-paying asset. In that same section, we also extend the Greeks to consider sensitivity of the option behavior to the contractual amortization rate. Section 4 then presents a brief discussion of applicability of this new option type for use in hedging for both traditional and decentralized finance settings. Section 5 concludes.
@@ 2 Amortizing Options
Installment options have been offered over-the-counter in a few markets: “installment warrants on Australian stocks listed on the Australian Stock Exchange (ASX) [10, 4], a 10-year warrant with nine annual payments offered by Deutsche Bank [7], and so on,”  [12, Section 1]. However, because lapsing destroys fungibility across units of notional, i.e. partial lapsing by making a partial installment payment is not permissible, traditional CI options are not amenable to exchange-based trading.
@@
Herein we propose a variant of CI options. Instead of directly paying the continuous installments, the option holder implicitly sells a fraction of his or her option holdings back to the underwriter at the current premium. Formally, if $N_{t}>0$ denotes the time $t$ claimable notional and $V_{t}>0$ denotes the option’s premium per unit notional, then funding the installment cost $c_{t}N_{t}dt$ with a sale of the notional leads to an exponential decay in the notional:
@@
Notably, because the notional is decreasing from these implicit payments, the requisite installment payments are also decreasing over time. In contrast to CI options in which the option underwriter is paid continuously in time, this option variant “pays” the underwriter via the decrease in his or her risk exposure.
@@
An amortizing perpetual option (AmPO) is an implicit payment, perpetual American CI option with installment costs $c_{t}=qV_{t}$ for some constant amortization rate $q>0$.
@@
While the fraction-of-premium installment payments ($c_{t}=qV_{t}$) implied by AmPO contracts is non-standard, it leads to a fixed, predictable amortization rate $q>0$ for the notional. In this way, the holder’s position deterministically decays over time, independent of the underlying market conditions. On the other hand, the underwriter is continuously compensated in proportion to the notional and fair-value exposure being carried; this holds because the implicit installment payments reduce that exposure in line with the contract’s value.
In contrast, under standard CI structures (as in Example 2.1), the amortization rate $c_{t}/V_{t}$ can spike or plummet with market conditions. For instance, if the underlying asset’s price remains constant, but its volatility rises, then $V_{t}$ increases while $c_{t}$ remains constant. Under such a scenario, the realized amortization rate would fall, leaving the underwriter with a disproportionately large remaining exposure.
@@
Alternative installment payments (e.g., as in Example 2.1) can be chosen instead, leading to time-varying amortization rates $c_{t}/V_{t}$. Regardless of the installment cost structure, amortizing CI variants become exchange tradable as every unit of notional is always identical and cannot lapse independently, i.e., all contract units are perfectly fungible. For such state-dependent amortization schedules, the valuation problem to be presented in Section 3.1 can be approached along the lines of [2]. For simplicity, herein we focus solely on the constant amortization case presented in Definition 2.2.
Remark 3.
Step options [15, 9] similarly feature payoffs with an effective notional that can decay over time. However, in contrast to the deterministic amortization of AmPOs, the decay in step options depends on the occupation time of the underlying, thus requiring path-dependent valuation and eliminating the fungibility required for exchange-based trading.

### v2

@@ Abstract
In this work, we introduce amortizing perpetual options (AmPOs), a fungible variant of continuous-installment options suitable for exchange-based trading. Traditional installment options lapse when holders cease their payments, destroying fungibility across units of notional. AmPOs replace explicit installment payments and the need for lapsing logic with an implicit payment scheme via the decay of the claimable notional. This amortization ensures all units evolve identically, preserving fungibility. We demonstrate that AmPO valuation can be reduced to an equivalent vanilla perpetual American option on a dividend-paying asset. In this way, analytical expressions are possible for the exercise boundaries and risk-neutral valuations for calls and puts. These formulas and relations allow us to derive the Greeks and study comparative statics with respect to the amortization rate. Illustrative numerical case studies demonstrate how the amortization rate shapes option behavior and reveal the resulting tradeoffs in the effective volatility sensitivity.
@@ 1 Introduction
Perpetual contracts are experiencing a renaissance due to their adoption in decentralized finance and cryptocurrency markets. Though introduced in [23] for, e.g., a market on real estate indices, perpetual futures on cryptocurrencies are liquidly traded on a number of markets such as Hyperliquid and Binance. These contracts readily allow traders to purchase leveraged positions or to short assets which, otherwise, are technologically incapable of such a position; we refer the interested reader to, e.g., [1, 3]. However, perpetual futures only provide users with linear payoffs. Perpetual options, which replicate elements of the funding rate structure of perpetual futures, have recently been introduced in decentralized systems. We highlight everlasting options [25, 1] and Panoptic [19] as specific perpetual option variants.
However, though perpetual futures have found wide usage in cryptocurrencies, perpetual options have not gained the same market share. Notably, the stochastic funding rates used in everlasting options and at Panoptic can be viewed as types of installment options. An installment option is one in which cash flows are exchanged between the long and short positions [12, 4]. Traditionally, these cash flows exist solely from the long to the short holders. Such options, however, only trade over-the-counter due to the nature of these contracts [4]. Specifically, installment options grant the holder a dual optionality: exercise or lapse. Because the lapsing logic comes from failure to make the installment payment in full, the notional units of these contracts are not identical and, therefore, cannot trade on an exchange. For the purposes of this work, we specifically highlight continuous-installment (CI) options which require a continuous stream of payments from the option holder to the option underwriter in order to keep the contract alive [10, 26, 17, 18, 8].
The remainder of this paper is organized as follows. In Section 2, we introduce an exchange-tradable variant of CI options; this is accomplished by replacing the explicit installment payments with implicit amortization of the options notional. This amortizing contract design eliminates the lapsing logic of CI options, thereby constructing a fungible option with a positive cost-of-carry; this structural reformulation is the primary contribution of this work. In Section 3, we study the risk-neutral valuation of these novel amortizing options by determining an equivalence with vanilla perpetual American options for a dividend-paying asset. In Section 3.2, under the Black-Scholes framework, we formulate the Greeks and also consider sensitivity of the option behavior to the contractual amortization rate. Section 4.1 provides considerations for how to select the amortization rate in practice. Section 4.2 then presents a brief discussion of applicability of this new option type in both traditional and decentralized finance settings. Section 5 concludes.
@@ 2 Amortizing Options
Throughout this work, we fix a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ satisfying the usual conditions. All stochastic processes introduced hereafter are assumed to be adapted to this filtration unless otherwise stated.
@@
Installment options have been offered over-the-counter in a few markets: “installment warrants on Australian stocks listed on the Australian Stock Exchange (ASX) [15, 4], a 10-year warrant with nine annual payments offered by Deutsche Bank [11], and so on,”  [17, Section 1]. However, because lapsing destroys fungibility across units of notional, i.e. partial lapsing by making a partial installment payment is not permissible, traditional CI options are not amenable to exchange-based trading.
@@
Herein we propose an exchange-tradable variant of CI options to maintain a positive cost-of-carry; as far as the author is aware, this is the first perpetual option design which is fungible with a positive cost-of-carry. Instead of directly paying the continuous installments, the option holder implicitly sells a fraction of his or her option holdings back to the underwriter at the current premium. Formally, if $N_{t}>0$ denotes the time $t$ claimable notional and $V_{t}>0$ denotes the option’s premium per unit notional, then funding the installment cost $c_{t}N_{t}dt$ with a sale of the notional leads to an exponential decay in the notional:
@@
Notably, because the notional is decreasing from these implicit payments, the requisite installment payments are also decreasing over time. In contrast to CI options in which the option underwriter is paid continuously in time, this option variant “pays” the underwriter via the decrease in his or her risk exposure. We note that the amortization rate $q_{t}$, if time $t$ dependent, is based on a global time-index and not based on the contract age; this construction is essential to maintain fungibility between contracts.
@@
An amortizing perpetual option (AmPO) is an implicit payment, perpetual American CI option with installment costs $c_{t}=q_{t}V_{t}$ for some progressively measurable amortization rate process $(q_{t})_{t\geq 0}$ such that $q_{t}\geq 0$ a.s. and $\int_{0}^{t}q_{s}ds<\infty$ a.s. for every time $t\geq 0$. That is, for an AmPO with payoff function $\Phi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ on underlying $(S_{t})_{t\geq 0}$ purchased at time $t_{0}$, the realized exercise value is $e^{-\int_{t_{0}}^{\tau}q_{s}ds}\Phi(S_{\tau})$ with exercise time $\tau$.
Proposition 2.3.
Consider an AmPO with amortization rate $(q_{t})_{t\geq 0}$ and payoff function $\Phi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ on some underlying $(S_{t})_{t\geq 0}$. A rational option holder, who purchased this AmPO at time $t_{0}\geq 0$, will never cancel the option position.
Proof.
Since the AmPO requires no explicit installment payments from the holder to maintain the position, the only cost-of-carry is the decay of the claimable notional. Note that this cost is internal to the valuation. Since the payoff function $\Phi(S)\geq 0$ for any price $S\geq 0$, the realized payout $e^{-\int_{t_{0}}^{\tau}q_{s}ds}\Phi(S_{\tau})\geq 0$ a.s. for any stopping time $\tau\geq t_{0}$. As this holds for any exercise time $\tau$, the value of the AmPO must also be non-negative. As the value of the cancelled AmPO is uniformly 0, the result follows. ∎
@@
Proposition 2.3 highlights a critical distinction between AmPOs and traditional CI options. In a traditional CI option, if the installment rate $c_{t}$ becomes large relative to the option value $V_{t}$ (e.g., as the option moves out-of-the-money), the option holder terminates the contract to stop the negative carry costs. In contrast, because AmPOs are, effectively, non-cancellable (as it is never optimal to cancel), as the option value $V_{t}$ drops, the effective amortization rate $q_{t}=c_{t}/V_{t}$ can explode (e.g., with the constant installment rate $c_{t}\equiv c>0$ as highlighted in Example 2.1). This jump in the amortization rate causes an accelerated liquidation of the notional. This creates an endogenous approximate lapsing of the option as the claimable notional decays to some minimal size. This effect is investigated further within Section 4.1.
Recall that AmPOs are amortizing options, meaning that the notional claim decays along the process $N_{t}=e^{-\int_{t_{0}}^{t}q_{s}ds}N_{t_{0}}$ for any times $t\geq t_{0}$. In this way, these perpetual options can be exercised at any stopping time $\tau$ resulting in a payoff of $e^{-\int_{t_{0}}^{\tau}q_{s}ds}\Phi(S_{\tau})$ for payoff function $\Phi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ per unit of notional claim with one unit of notional purchased at time $t_{0}$. However, option holders often want to maintain a constant notional exposure. To do so, the implicit installments paid through amortization become realized explicit installments paid to the market. That is, the option holder would pay $q_{t}V_{t}dt$ per unit of notional exposure to maintain the desired position. Because AmPOs are fungible, these payments are not necessarily made to the original underwriter who is compensated by the amortization. In Section 3.1, we will demonstrate that the valuation of these two viewpoints (decaying exposure with implicit payments or constant exposure with explicit option purchases) are equivalent problems so long as the amortization rate is an exogenous process.
@@
Step options [20, 13] similarly feature payoffs with an effective notional that can decay over time. However, in contrast to the amortization schedule of AmPOs, the decay in step options depends on the occupation time of the underlying, thus requiring path-dependent valuation and eliminating the fungibility required for exchange-based trading.
@@ 3 Risk-Neutral Valuation
Within this section, we want to investigate the risk-neutral valuation of these AmPOs as presented in the prior section. Throughout this section we will consider a market with risk-free rate $(r_{t})_{t\geq 0}$ and risky price process $(S_{t})_{t\geq 0}$. We, further, assume throughout this section that this market admits no arbitrage and, thus, there exists a risk-neutral measure $\mathbb{Q}\sim\mathbb{P}$ by the fundamental theorem of asset pricing. Within this system, we consider the risk-neutral valuation of a single claimable unit of notional of an AmPO with payoff function $\Phi:\mathbb{R}_{+}\to\mathbb{R}_{+}$. For simplicity of exposition, we formulate the valuation problem at time $t=0$; the general valuation at time $t>0$ follows comparably from a time-shift of the relevant processes. In this way, throughout this section we consider the optimal exercise problem
$V_{0}=\sup_{\tau\in\mathcal{T}}\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{0}^{\tau}(r_{s}+q_{s})ds}\Phi(S_{\tau})\right]$ (3.1)
where $\mathcal{T}$ denotes the set of all bounded stopping times with respect to the market filtration.
Assumption 3.1.
For the remainder of this work, we consider the payoff function $\Phi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ such that $\Phi(S_{\tau})\in L^{1}$ for all $\tau\in\mathcal{T}$. In addition, we assume the interest rate $(r_{t})_{t\geq 0}$ and amortization $(q_{t})_{t\geq 0}$ processes are progressively measurable such that $r_{t},q_{t}\geq 0$ a.s. and $\int_{0}^{t}r_{s}ds,\int_{0}^{t}q_{s}ds<\infty$ a.s. for every time $t\geq 0$.
In studying this valuation problem, we consider two settings. First, in Section 3.1, we consider the general valuation problem by determining an isomorphism between AmPOs and perpetual American options on dividend-paying assets. Then, in Section 3.2, we specifically investigate the Black-Scholes framework for AmPOs with constant amortization rates $q_{t}\equiv q>0$ to provide intuition on the pricing and hedging of these instruments.
3.1 General Setting
To determine the risk-neutral value of an AmPO, we first observe a fundamental symmetry in the value process. Holding an AmPO involves a decaying notional $N_{t}=e^{-\int_{0}^{t}q_{s}ds}N_{0}$ on the options payoff $\Phi(\cdot)$. That is, from the holder’s perspective, the position’s value is $N_{t}V_{t}$ for AmPO option value $(V_{t})_{t\geq 0}$ with payoff $N_{\tau}\Phi(S_{\tau})$ at exercise time $\tau$. The amortization immediately shows up as an additional discounting term (i.e., an updated effective risk-free rate $(r_{t}+q_{t})_{t\geq 0}$); with this augmented risk-free rate, the amortization also functionally appears as a drag on the modified drift to the underlying asset (i.e., an effective dividend yield augmented by $(q_{t})_{t\geq 0}$). In this way, in Lemma 3.2, we demonstrate that AmPO valuation perfectly coincides with (vanilla) perpetual American options under modified parameters.
Lemma 3.2.
Let the risky asset price process follow $(S_{t})_{t\geq 0}$ with dividend rate $(\delta_{t})_{t\geq 0}$. The AmPO with amortization rate $(q_{t})_{t\geq 0}$ is priced identically to the perpetual American option with the same payoff, effective risk-free rate $(r_{t}+q_{t})_{t\geq 0}$, and dividend yield $(\delta_{t}+q_{t})_{t\geq 0}$.
Proof.
Consider the money-market account $B_{t}:=\exp(\int_{0}^{t}r_{s}ds)$ for any time $t$ and assume the risky asset pays a (possibly time-varying) dividend yield $(\delta_{t})_{t\geq 0}$ in the sense that the discounted, dividend-reinvested price process is a $\mathbb{Q}$-local martingale, i.e., $(B_{t}^{-1}\exp(\int_{0}^{t}\delta_{s}ds)S_{t})_{t\geq 0}$ is a $\mathbb{Q}$-local martingale. Furthermore, define the effective money-market account $\tilde{B}_{t}:=\exp(\int_{0}^{t}(r_{s}+q_{s})ds)$. Immeidately, as $\tilde{B}_{t}^{-1}\exp(\int_{0}^{t}(\delta_{s}+q_{s})ds)S_{t}=B_{t}^{-1}\exp(\int_{0}^{t}\delta_{s}ds)S_{t}$, we note that $S$ with effective dividend yield $(\delta_{t}+q_{t})_{t\geq 0}$ is a $\mathbb{Q}$-local martingale with discounting by $\tilde{B}$ . Now consider the AmPO value at time $t=0$ provided in (3.1), i.e., $V_{0}=\sup_{\tau\in\mathcal{T}}\mathbb{E}^{\mathbb{Q}}[\tilde{B}_{\tau}^{-1}\Phi(S_{\tau})]$. By observation, the stopping problem defining $V_{0}$ is exactly the perpetual American option valuation problem in a market with risk-free rate $(r_{t}+q_{t})_{t\geq 0}$ and dividend yield $(\delta_{t}+q_{t})_{t\geq 0}$. ∎
Lemma 3.2 provides a direct methodology to study the risk-neutral valuation of AmPOs in general settings. By solving the equivalent perpetual American option valuation problem, the valuation and exercise boundary for AmPOs can be determined. We refer the interested reader to, e.g., [7, 6] for theory on pricing these options under general Lévy processes and to [2] for a methodology for such problems under jump-diffusion processes. In Section 3.2 below, we will detail pricing and hedging in the Black-Scholes framework under a constant amortization rate $q_{t}\equiv q>0$ to provide a simple setting with closed form solutions.
Though we found that AmPOs can be priced as perpetual American options on dividend-paying assets with modified market parameters, we note that these still function as CI variants. In Proposition 3.3, we demonstrate that the AmPO valuation problem is identical to that of the CI option with $c_{t}=q_{t}V_{t}$ (assuming constant claimable notional) if the amortization rate is exogenously defined from the option value. Such results are akin to those found in pricing credit risk [14].
Proposition 3.3.
Fix an amortization rate process $(q_{t})_{t\geq 0}$. Define the operator $\mathcal{C}$ on the class of non-negative adapted càdlàg processes $V$ such that $\int_{0}^{t}q_{s}V_{s}ds\in L^{1}$ for all $t\geq 0$ by
$\mathcal{C}(V)_{t}:=\operatorname*{ess\,sup}_{\tau\in\mathcal{T},\tau\geq t}\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t}^{\tau}r_{s}ds}\Phi(S_{\tau})-\int_{t}^{\tau}e^{-\int_{t}^{s}r_{u}du}q_{s}V_{s}ds\;\bigg|\;\mathcal{F}_{t}\right]\vee 0.$
The AmPO value process $(V_{t})_{t\geq 0}$ satisfies the fixed point $V=\mathcal{C}(V)$. That is, $V$ is a valid price process for the self-referential proportional-installment CI contract. Furthermore, the lapse feature of this CI contract is never optimal.
Proof.
Define the conditional payoff mapping $\bar{\mathcal{C}}^{\tau}$ for a stopping time $\tau\in\mathcal{T}$ such that
$\bar{\mathcal{C}}^{\tau}(V)_{t}:=\mathbb{E}^{\mathbb{Q}}\left[e^{-\int_{t\wedge\tau}^{\tau}r_{s}ds}\Phi(S_{\tau})-\int_{t\wedge\tau}^{\tau}e^{-\int_{t\wedge\tau}^{s}r_{u}du}q_{s}V_{s}ds\;\big|\;\mathcal{F}_{t\wedge\tau}\right]$
for any non-negative adapted càdlàg process $V$. Let $V=(V_{t})_{t\geq 0}$ be the AmPO value process. Define the discount factor $D_{t}:=\exp(-\int_{0}^{t}r_{s}ds)$ and the cumulative amortization factor $Q_{t}:=\exp(-\int_{0}^{t}q_{s}ds)$. From the definition of the American option value in Lemma 3.2, the amortized, discounted value process $M_{t}:=D_{t}Q_{t}V_{t}$ is a $\mathbb{Q}$-supermartingale for all $t$. Furthermore, the stopped process $(M_{t\wedge\tau^{*}})_{t\geq 0}$ is a $\mathbb{Q}$-martingale, where $\tau^{*}\in\mathcal{T}$ is the optimal exercise time for the AmPO.
Consider $\bar{\mathcal{C}}^{\tau}(V)_{t}$ for an arbitrary stopping time $\tau\in\mathcal{T}$ (assuming $\tau\geq t$ a.s. without loss of generality):
$\bar{\mathcal{C}}^{\tau}(V)_{t}$ $=\frac{1}{D_{t}}\mathbb{E}^{\mathbb{Q}}\left[D_{\tau}\Phi(S_{\tau})-\int_{t}^{\tau}D_{s}q_{s}V_{s}ds\,\bigg|\,\mathcal{F}_{t}\right]$
$=\frac{1}{D_{t}}\mathbb{E}^{\mathbb{Q}}\left[D_{\tau}\Phi(S_{\tau})-\int_{t}^{\tau}q_{s}Q_{s}^{-1}M_{s}ds\,\bigg|\,\mathcal{F}_{t}\right]$
$=\frac{1}{D_{t}}\mathbb{E}^{\mathbb{Q}}\left[D_{\tau}\Phi(S_{\tau})-\int_{t}^{\tau}M_{s}d(Q_{s}^{-1})\,\bigg|\,\mathcal{F}_{t}\right]$
$=\frac{1}{D_{t}}\mathbb{E}^{\mathbb{Q}}\left[D_{\tau}\Phi(S_{\tau})-Q_{\tau}^{-1}M_{\tau}+Q_{t}^{-1}M_{t}+\int_{t}^{\tau}Q_{s}^{-1}dM_{s}\,\bigg|\,\mathcal{F}_{t}\right].$
Recall that $Q_{\tau}^{-1}M_{\tau}=D_{\tau}V_{\tau}$ and $Q_{t}^{-1}M_{t}=D_{t}V_{t}$. Substituting these back into the expectation:
$\bar{\mathcal{C}}^{\tau}(V)_{t}=V_{t}+\frac{1}{D_{t}}\underbrace{\mathbb{E}^{\mathbb{Q}}\left[D_{\tau}(\Phi(S_{\tau})-V_{\tau})\,\bigg|\,\mathcal{F}_{t}\right]}_{\leq 0}+\frac{1}{D_{t}}\underbrace{\mathbb{E}^{\mathbb{Q}}\left[\int_{t}^{\tau}Q_{s}^{-1}dM_{s}\,\bigg|\,\mathcal{F}_{t}\right]}_{\leq 0}\leq V_{t}.$
The first term is non-positive because the American option value always exceeds the immediate payoff ($V_{\tau}\geq\Phi(S_{\tau})$). The second term is non-positive because $\int Q^{-1}dM$ is a $\mathbb{Q}$-local supermartingale (as $M$ is a $\mathbb{Q}$-supermartingale and $Q^{-1}$ is positive). However, for the specific choice $\tau=\tau^{*}$, $V_{\tau^{*}}=\Phi(S_{\tau^{*}})$ (by optimal exercise) and $M$ is a $\mathbb{Q}$-martingale (up to $\tau^{*}$), so both terms vanish. Thus, $\bar{\mathcal{C}}^{\tau^{*}}(V)_{t}=V_{t}$. Taking the supremum over $\tau$, we find $\mathcal{C}(V)_{t}=V_{t}\geq 0$, completing the proof. ∎
Notably, the requirement in Proposition 3.3 of an exogenously defined amortization rate process $(q_{t})_{t\geq 0}$ (i.e., one that is not a direct function of $V$) is not satisfied in standard CI option constructions (see Example 2.1). In the following section, we will consider the specific case of a constant amortization rate $q_{t}\equiv q>0$; this simplifies the setup for explicit valuation while trivially satisfying the regularity conditions imposed in Proposition 3.3.
3.2 Constant Amortization in the Black-Scholes Framework